2b+3a+4a=
b=6
a=5
what's the answer Please....

Answer by ewatrrr(24785)


Hi,
*Note: the idea is to subsitute the values given for a and b
a = 5, b = 6
2b + 3a + 4a=
2*6 + 3*5 + 4*5 =
Think You can handle it from here as to showing Your work
47 is what it adds up to